OpenStudy (anonymous):

if a bicycle has 26-inch diameter wheels (the front chain drive has a radius of 4 inches and the backdrive has a radius of 1 inch), how far does the bicycle travel every one rotation of the pedals? round to the nearest inch.

OpenStudy (anonymous):

The circumference of a circle is pi times the diameter. Diameter is twice the radius. \[\text{Round}\left[\frac{(8 \pi ) 26 \pi }{2 \pi },1\right]=327 \]